OpenStudy (anonymous):

Calculate the derivative of the following function. g(x)=sin(x^2)+(sinx)^2+sin(2x)+(sin2)x

myininaya (myininaya):

\[y=f(g(x)) =>y'=f'(g(x))g'(x)\] use chain rule

OpenStudy (anonymous):

could you write the process?

OpenStudy (anonymous):

you have to take the derivative of 4 terms. we already did the last one

OpenStudy (anonymous):

derivative of \[\frac{d}{dx}\sin(x^2)=\cos(x^2)\times \frac{d}{dx}x^2=2x\cos(x^2)\] for the first term

OpenStudy (anonymous):

g'(x)=2 cos(2 x)+2 x cos(x^2)+sin(2)+sin(2 x)